Leonardo, who is married but files separately, earns $80,000 of taxable income. He also has $15,000 in city of Tulsa bonds. His wife, Theresa, earns $50,000 of taxable income.

What is Leonardo and Theresa's effective tax rate for year 2018?
Answer #1

Answer) 14.12%

Explanations:

Total Income = 145000

Taxable Income = 145000 - 15000 = 130000

Tax = 8907 + ( 130000 - 77400) x 22% = 20479

Effective tax rate = 20479 / 145000 = 14.12%